Recent Log Entries

CREW CHIEF SIGN IN

Log Number: 4104985Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 01-05-23
Time : 07:12:10
Day : Thursday
Shift : Day

Crew Chief : Humphry
Program Deputy: Spata
Team : X
Arm : RadCon
Designated SSO: Christian
LERF Operator:
Operators : Christian

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

Hall A Laser PMES Unstable

Log Number: 4104982Logbook(s): Entry Makers: deirBody: 

EGG took a look at it on swing shift and gave the okay to adjust alarm limits, ~21:00.

GMES has been going up and down ever since. Slightly adjusted the LOLO to keep it from alarming.

Fig 3&4: Looks like it could be related to phase fluctuations.

Operator SIGN IN

Log Number: 4104975Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Deir
DATE:             01/05/2023
SHIFT:               Owl

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

Injector work summary

Log Number: 4104969Logbook(s): Entry Makers: ywanghoflerBody: 

Optimized beam up to Fcup # with HWien angle at -64.59 deg.
Established high current to Fcup #1:
Fig. 1---45uA, Laser A
Fig. 2---65uA, Laser C
Fig. 3---110uA ( A+ C)

Autolog save

Log Number: 4104968Logbook(s): Tags: Body: saver ran from opsl06

Saved: ywang 01-04-23 21:47:06 {A:3 B:5 C:5 E:1047 A=CW B=BS C=BS D=BS 0R08=0.00 0R08 dump in. dump(s) in.

Autolog save

Log Number: 4104963Logbook(s): Tags: Body: saver ran from opsl06

Saved: ywang 01-04-23 20:49:17 {A:3 B:5 C:5 E:1047 A=BS B=BS C=CW D=BS 0R08=0.00 0R08 dump in. dump(s) in.

COMMENT ON ARC3 PS Mains High Power Testing Into Resistive Load and Commissioning

Log Number: 4104948Logbook(s): Body: 

Mark and I ran repaired ARC3 power supply on a resistive load (2.33 Ohms) today up to 450A. Looking at the diagnostic waveform on NSH side @ 300A and 450A we see that the current sharing among all four NSH4 sections aren't equal rather NSH3 is running around 10% higher than rest of them. Please note NSH3 converter section is where we replaced the bad choke last year.

Tomorrow we will look into the PWM board because if the reference triangular voltage is higher then it will drive the IGBT harder and run higher current depending on the percentage off.

P.S.

QE Measurement

Log Number: 4104944Logbook(s): Body: Spot Location: (10700.0,18500.0), Charge: 4.2000000000000004e-5

Laser Wavelength Atten Setting Laser Power Current HV PS Current QE PV Updated? Update Time/Reason Not Updated (nm) (W) (uA)</ce

1L12-2 Cavity Frequency after Cryo Cooldown

Log Number: 4104939Logbook(s): Entry Makers: xgomezgelhaarBody: 

Xavier and I verified cavity frequency was at 1497.100000Mhz after cooldown was complete. Measured at 4 and 5 pm. Stepper tuner program running and and cavity will be checked in the morning before stopping it and disabling tuner.

QE Measurement

Log Number: 4104938Logbook(s): Body: *New Spot* Spot Location: (10700.0,18500.0)

Laser Wavelength Atten Setting Laser Power Current HV PS Current QE PV Updated? Update Time/Reason Not Updated (nm) (W) (uA)

Text alert added to 10Amp trim alarms

Log Number: 4104930Logbook(s): Body: 

Received a request to add a text alert if a 10Amp trim goes into MAJOR alarm.
Alerts will be sent to Jim Colman and Kevin Banks.
Restarted the alarm handlers at with the permission of the crew.

Autolog save

Log Number: 4104925Logbook(s): Tags: Body: saver ran from opsl06

Saved: ywang 01-04-23 16:37:50 {A:3 B:5 C:5 E:1047 A=Tune B=Pulse C=Pulse D=BS 0R08=0.00 0R08 dump in. dump(s) in.

Pages